OpenStudy (anonymous):

2x>-6 do you add 2 to 6?

OpenStudy (anonymous):

No, you divide -6 by 2

OpenStudy (anonymous):

x>-3

OpenStudy (anonymous):

Correct :)

OpenStudy (saifoo.khan):

\[\checkmark\]

OpenStudy (anonymous):

how do i graph this?

OpenStudy (saifoo.khan):



OpenStudy (saifoo.khan):

As you can see, "x" is greater than -3.

OpenStudy (mysesshou):

And also not including -3 because of the circle. :)

OpenStudy (anonymous):

negative a open circle and postive a closed one

OpenStudy (saifoo.khan):

Oh right.

OpenStudy (anonymous):

i got this one x-4<(with line under it) 3. im graphing those two on same line

OpenStudy (mysesshou):

"greater than equal to " or "equals" is a closed circle, including the number at the circle. "greater than" means open circle because it is only up to the number not including

OpenStudy (anonymous):

> is greater than or less than?

OpenStudy (mysesshou):

depends where the numbers and x are. for x > -3 it is read x is greater than -3. Imagine it like an alligator mouth. it wants to eat the most, so the x is the greater than.

OpenStudy (mysesshou):

if it was written x < -3 then it is "x is less than -3" because the alligator is going for the -3 then it is read as the opposite for the "x is ____"

OpenStudy (mysesshou):

not sure i understand this Q: i got this one x-4<(with line under it) 3. im graphing those two on same line

OpenStudy (mysesshou):

oooohh.

OpenStudy (mysesshou):

\[x-4 \le 3\] ?

OpenStudy (anonymous):

yeaa

OpenStudy (mysesshou):

ok. add 4 to both sides.

OpenStudy (mysesshou):

\[x \le 7\]

OpenStudy (mysesshou):

you're adding this to the previous graph?

OpenStudy (anonymous):

yes

OpenStudy (mysesshou):

\[x >-3\] and \[x \le 7\] what you do is open circle for -3 because it is not including it. filled in circle for the 7 because it is including it. you can rewrite it to \[-3<x\]and \[x \le 7\] and then combine it \[-3<x \le 7\] so draw the line. then put a open circle on -3 and a filled in circle on 7. then connect a line between them. ok?

OpenStudy (mysesshou):